I need help with a linear word problem and equation (Please explain how you got your answer because I'm confused. Thank you.):
Fundrasing: To raise money, your student council is selling magazine subscriptions. The student council will receive a one-time bonus of $150 from the magazine publisher plus 38% of the subscription money. The following verbal model represents the situation. Write a linear equation from the verbal model.
Money = Bonus + Subscription . Subscription
Raised Percentage Money
What I got was m=150+0.38x.
One more thing:
How much subscription money is needed for the council to raise a total of $300? Round your answer to the nearest dollor.

Found 2 solutions by solver91311, mananth:
Answer by solver91311(24713)



Yep,



is the linear relationship where represents the money raised, and is the amount of income from subscriptions.

If you want , then solve



for

You might want to round your answer to the nearest "dollar." I don't know what a "dollor" is.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by mananth(16946)


To raise money, your student council is selling magazine subscriptions. The student council will receive a one-time bonus of $150 from the magazine publisher plus 38% of the subscription money. The following verbal model represents the situation. Write a linear equation from the verbal model.

let the subscription be x
they recieve 38% of the
per subscription they receive 0.38x
M = 150+0.38x
300=150+0.38x
-150
300-150=0.38x
150=0.38x
/0.38
150/0.38=x
$3947 is the subscription money required
...
m.ananth@hotmail.ca